OpenStudy (stormswan):

Solve for x: −3|2x + 6| = −12 I'm confused on how to solve this. I am an exchange student, and I haven't relly learn this stuff yet.

OpenStudy (mathmale):

Welcome to the USA, and to OpenStudy. From which country are you? −3|2x + 6| = −12 involves "absolute value." I'd suggest you look that up, either on the Internet or in your textbook (if you h ave one) or course materials (if you have any).

OpenStudy (mathmale):

Here's one way to explain what the absolute value operator does, in context: (1) if 2x+6 is already positive, |2x+6| = 2x+6. (2) if 2x+6 happens to be negative, then |2x+6|=-(2x+6). So you end up with two separate equations to solve. Try this. Feel free to ask questions about anything that remains unclear to you. Delighted that you've chosen to post this question.

OpenStudy (nikato):

First thing you should do is divide -3 on both sides

OpenStudy (mathmale):

Nikato's suggestion would definitely help to simplify the given equation (but is not essential to solving the problem).

OpenStudy (stormswan):

so the answer to nikato 's post would be 6 right?

OpenStudy (stormswan):

so like |2x + 6| = 6 ?

OpenStudy (stormswan):

@mathmale

OpenStudy (mathmale):

There will be TWO solutions. -3|2x + 6| = −12 reduces to (a) 2x+6=4 => 2x=-2. Find x. (b) -(2x+6)=4. Find this (different) value of x. The results are your solution set.

OpenStudy (stormswan):

too solution (a) x would be -1?

OpenStudy (mathmale):

Yes. Remember, you don't need me to check your solutions; all you have to do to check them is to substitute them, one at a time, back into the original equation. If the orig. eqn. is true, your x is a solution; otherwise, it is not a solution.

OpenStudy (stormswan):

ok, i think i get it. so it is x= -1 and x= -5? just cheking

OpenStudy (mathmale):

SS: I think it's important that YOU check in the manner I've suggested, rather than asking me or others whether you are "right" or not. I don't mean to be unkind.

OpenStudy (stormswan):

ok, i understand. tank you so very much.